OpenStudy (anonymous):

the triceta at the race car consists selecting of first second and 3rd finishers in a pirpicular order. if they are elevl entries in the race, how many pruchaces must you purchace to garatee a win

OpenStudy (anonymous):

You want a permutation of 11 choosing 3: 11p3 = 11!/(11-3)! = 11*10*9 = 990

OpenStudy (anonymous):

990 tickets?

OpenStudy (anonymous):

yes.